26' A vehicle starting its motion from rest travels at a uniform acceleration of 5ms-2 for a certain time and



travels with the attained uniform velocity for another certain period of time. Then it decelerates uniformly at



5ms-2 and stops after a certain time. If the total time spent is 25s and the average velocity for the whole



motion is 20ms-1 , find the time that the vehicle travelled at uniform velocity.

1
Expert's answer
2022-02-06T14:33:50-0500

From the data given, we understand that the average velocity equals half the velocity of uniform motion:


"v_\text{av}=at/2,\\\space\\
t=2v_\text{av}/a=10\text{ s}."

The time that the vehicle travelled at uniform velocity:


"t_\text u=2T-t=25-2·10=5\text{ s}."